Fermi National Laboratory


Accelerator Update

Wednesday July 21
The day shift began with Operations monitoring store 3663 and the Recycler stash, but with no Antiproton Source (Pbar) stacking or beam to MiniBooNE due to trouble with the I- Source.

A Preaccelerator expert replaced the brushes for the I- Source motor generator and then Pbar resumed stacking at 8:52 AM.

At 9:55 AM, a Linac RF station (KRF2) had tripped off twice in five minutes. It reset okay.

At 10:11 AM, a Linac RF station (LRF5) had tripped off twice in one minute. It reset okay.

At 10:24 AM, a Booster RF station (BRF9) tripped off. It reset okay.

At 10:57 AM, a Booster RF station (BRF1) tripped off. It reset okay.

At 11:18 AM, KRF2 tripped off again. Linac technicians helped Operations reset okay.

At 11:56 AM, a Booster power supply (GMPS) and a RF station (BRF11) tripped off. Both reset okay.

A Main Injector (MI) expert began a slipstacking study.

At 1:56 PM, a Tevatron (TeV) corrector power supply (T:CV49) tripped off due to a blown fuse. At the same time colliding beam problems appeared for the Collider Detector Facility at Fermilab (CDF) and D-Zero (D0). An expert investigated.

At 3:58 PM, KRF5 had tripped off twice in three minutes. It reset okay.

At 4:34 PM, the Tevatron Electron Lens (TEL) tripped off on phase imbalance. It reset okay.

At 5:12 PM, Operations reported that Pbar started to unstack. Operators investigated.

Operations terminated store 3663 at 6:55 PM.

Operations put the TeV through a dry squeeze at 7:49 PM and then into shot setup at 8:23 PM.

Operations established store 3665 at 10:32 PM with an initial luminosity of 86.48E30. Pbar resumed stacking.

A Pbar expert began a Debuncher Accumulator study at 11:02 PM.

Thursday July 22
The midnight shift began with Operations monitoring store 3665, with Pbar conducting studies, and with MiniBooNE taking beam.

Pbar studies ended at 12:52 AM.

Operations reported at 1:03 AM that they were having problems establishing beam to Switchyard (SY) due to SEPTA positions. This was affecting stacking, so operators disabled beam to SY.

At 5:49 AM, Operations reported that a Linac quadrupole power supply (QPS133) was drifting high and causing beam trouble. It came back into normal parameters.

At 7:31 AM, Operations turned off beam to Linac, Booster, MI, Pbar, and MiniBooNE in preparation for the four-hour morning access.

By 12:40 PM, Operations had Linac and Booster back up with Booster expert conducting GMPS studies (no beam).

Operations established beam to Pbar at 1:46 PM.

Operations was running low intensity beam to the Booster for tuning. A Booster expert reported that the notcher wasn't firing. Booster techs investigated.

Operations again established beam to Pbar at 2:55 PM, but the MI low level RF system developed a problem.

Pbar began stacking at 4:49 PM.

A Linac expert responded to a flurry of Linac quadrupole (QPS133) trips.

Operations established beam to MiniBooNE at 5:28 PM.

After talking to a Linac expert, operators replaced all the control cards for QPS133 at 6:39 PM. This seemed to help.

At 7:36 PM, a 480-volt breaker tripped at the West Utility yard. This caused many RF stations to trip off in Linac and Booster.

Operations established beam to Pbar and MiniBooNE at 8:30 PM.

Operations had established beam to the Meson MTest and MCenter experiments by 9:42 PM.

A Linac tech came in at 10:46 PM to investigate more QPS133 trips.

Operations reported at 11:47 PM that they had lost 80mA of antiprotons from the stack. Operators investigated.

Friday July 23
The midnight shift began with Operations monitoring store 3665, with Pbar Debuncher studies, and with beam to MiniBooNE and the MTest and MCenter experiments.

Pbar resumed stacking at 12:54 AM. However, an Accumulator SEPTA (ISEPV) was occasionally tripping off.

At 2:45 AM, Operations reported that they were still having problems with QPS133.

At 3:34 AM, the TeV suffered a sector E1 quench due to orbit distortions.

TeV experts began a study period at 6:30 AM.

Plans for Friday Morning and the Weekend
The plan for today is to allow machine experts to conduct studies for four hours or so. Late this afternoon Operations will establish a store. The weekend will consist of stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/23/2004   email Fermilab